I messed up on my last question, can you please show how to solve this quadratic equation by completing the square not sq. rt prop.
m^2 = 8m + 3

Answer by nerdybill(7384)

m^2 = 8m + 3
Move all the 'm' terms to one side:
m^2 - 8m = 3
Take half of b (-8) and square it (16). Add to both sides:
m^2 - 8m + 16 = 3 + 16
Factoring the left side:
(m-4)(m-4) = 19
(m-4)^2 = 19
Take the square root of both sides
m-4 = sqrt(19)
m = 4+sqrt(19)
or
m = {8.3589, -0.3589}